Number of carbon atoms in 2.5g of HC2H3O2

1
Expert's answer
2022-02-20T08:39:41-0500

RMM of "HC_2H_3O_2=1+2\times 12+1\times3+2\times 16=60gm"

Number of moles of "HC_2H_3O_2" ="2.5\/60=0.04167"

Number of carbon atoms in 2.5g of "HC_2H_3O_2" ="0.04167\times 2\times 6.022\times 10^{23}=5.0187\times 10^{22}"